# Fabien Despinoy

### v2

@@ 4.2 Annotation-Free Semantic Segmentation Results
Table 1 reports the mIoU on the different datasets. Results of methods in first block are taken from TCL [Cha et al.(2023)Cha, Mun, and Roh] while second block methods are taken from their corresponding papers.
@@
First, training free methods, CLIP-DIY [Wysoczańska et al.(2024)Wysoczańska, Ramamonjisoa, Trzciński, and Siméoni], SCCLIP [Wang et al.(2023a)Wang, Mei, and Yuille], CaR [Sun et al.(2023)Sun, Li, Torr, Gu, and Li] and ReCo [Shin et al.(2022)Shin, Xie, and Albanie] perform inferior to our method mostly with a large margin except for CaR [Sun et al.(2023)Sun, Li, Torr, Gu, and Li] on PV-21 which can be attributed to the post-processing of background in CaR. Second, FMbSeg performs best or second best on all datasets even improving over TCL [Cha et al.(2023)Cha, Mun, and Roh] trained on 15M images with image level labels. FMbSeg outperforms TCL by a large margin of 8% on PC-59 and 26% on CO-80. Third, FMbSeg outperforms SAM-CLIP [Wang et al.(2023b)Wang, Vasu, Faghri, Vemulapalli, Farajtabar, Mehta, Rastegari, Tuzel, and Pouransari] significantly on PV-21 and PC-59 while being inferior on CS-171 and ADE albeit with a small margin. SAM-CLIP [Wang et al.(2023b)Wang, Vasu, Faghri, Vemulapalli, Farajtabar, Mehta, Rastegari, Tuzel, and Pouransari] trains SAM encoder on 41M images with no code or model available which makes it infeasible for us to evaluate it on the remaining datasets. Nevertheless, we emphasize that our alignment module is light and trained only for a few epochs on almost 400 times less data compared to  [Wang et al.(2023b)Wang, Vasu, Faghri, Vemulapalli, Farajtabar, Mehta, Rastegari, Tuzel, and Pouransari]. This makes our method a plug and play approach to semantically segment any dataset.
Due to space limits in the main paper, we extensively ablate different ways of leveraging our approach in Appendix. We evaluate stage 1.1 and 1.2 performance as training free methods. We also consider a semi supervised variant where image level labels are provided to stage 1.1 which significantly improves the performance leading to SOTA performance without any training. In Appendix we further examine our alignment module (pretrained on COCO-Stuff) when fine-tuned on specific target datasets mentioned in table 1 with no ground-truth annotations and in a self-supervised manner. This fine-tuning does not add a large over-head due to the efficiency of our lightweight alignment module. As shown in Appendix, in this case, the performance on ADE20k increases to 21.57%.
@@ 4.3 Qualitative Results
First we inspect the alignment of patch features with vocabulary after we train our alignment module. Fig. 2 shows the patch level similarity between the image features and the image level text features on Pascal VOC. Notice that the newly aligned image encoder produces distinctly more consistent similarity heatmaps for the given categories than those produced by CLIP. This shows the efficacy of aligning a better vision encoder with text rather than trying to improve CLIP patches features which typically requires large scale (re)training.
Next, Fig. 4 shows qualitative results of the annotation-free zeroshot segmentation by FMbSeg on the different datasets.
@@ 5 Conclusion and Future Work
Given the impressive performance of vision-language foundation models and their transferability to various downstream tasks, we consider the problem of open vocabulary annotation-free semantic segmentation. By composition of different foundation models, namely CLIP and SAM, we extract free pixel level annotations. We then propose a lightweight alignment module that projects the embedding of any arbitrary pretrained vision encoder to the text encoder space. By pretraining on COCO-Stuff, we obtain a generalist alignment module that can perform semantic segmentation across a wide range of object categories. Alternatively, our method can be deployed as a plug-and-play customized alignment module for any semantic segmentation dataset with zero annotations. We show SOTA results and open a new direction for zeroshot semantic segmentation by foundational models compositionality. As future work, we want to investigate other image encoders and continuous fine-tuning on new categories. Additionally, we want to explore how our alignment module can improve VLM models that are based on CLIP to further strengthen their object localization capabilities.

### v3

@@ 4.2 Annotation-Free Semantic Segmentation
Table 1 reports the mIoU on different datasets. Results of methods in first block are taken from TCL [8] while second block methods are taken from their corresponding papers. The third block reports our Stage 2 results when trained with the pseudo annotations generated on COCO-Stuff dataset.
@@
First, training free methods, CLIP-DIY [34], SCCLIP [32], CaR [31] and ReCo [30] perform inferior to our method mostly with a large margin. Second, our method performs the best or second best on all datasets even improving over TCL [8] trained on 15M images with image level labels. FMbSeg-Stage 2 (base) outperforms TCL by a large margin of 8% on PC-59 and 26% on CO-80. Third, FMbSeg-Stage 2 (base) outperforms SAM-CLIP [33] significantly on PV-21 and PC-59 while being inferior on CS-171 and ADE albeit with a small margin. SAM-CLIP [33] trains SAM encoder on 41M images with no code or model available which makes it infeasible for us to evaluate it on the remaining datasets. Nevertheless, we emphasize that our Stage 1 method is training free and our alignment module is light and trained only for a few epochs on almost 400 times less data compared to SAM-CLIP [33]. This makes our method a plug and play approach to semantically segment any dataset. Finally, adding refinement to FMbSeg pixel segmentation brings an average improvement of $1.7\%$.
@@ 4.3 Qualitative Results
First we inspect the alignment of patch features with vocabulary after we train our alignment module. Fig. 2 shows the patch level similarity between the image features and the image level text features on Pascal VOC. The newly aligned image encoder produces distinctly more consistent similarity heatmaps for the given categories than those produced by CLIP. While the most lit patches in CLIP’s heatmaps typically correspond to the queried object, CLIP’s patch features are not sufficient to generate semantic segmentation masks of the objects. This shows the efficacy of aligning a better vision encoder with text rather than trying to improve CLIP patches features which typically requires large scale (re)training. Fig. 4 shows qualitative results of the annotation-free zeroshot segmentation by FMbSeg-Stage 2 (refined) on different datasets.
4.4 Ablation
4.4.1 Stage 1 Ablation
Table 2 evaluates the performance for the training-free part of our method, FMbSeg-Stage 1. The generated pseudo annotations for each dataset are evaluated against the dataset’s groundtruth segmentation masks. For ablation purposes only, we further provide the results for a semi-supervised variation of Stage 1.1 when the object presence detection part of the method (see 3.2.1) is replaced with the groundtruth image level annotations from the dataset.
Stage 1.1 performs the lowest due the limitations mentioned earlier, namely, small objects, single instance segmentations and wrong detections due to visual/textual ambiguities (fig 3). Stage 1.2 achieves a relatively better performance compared to Stage 1.1 as it addresses the limitations with small objects and can segment multiple instances. Semi-supervised Stage 1.1 removes the wrongly detected objects from the pipeline to vastly improve the performance, achieving comparable results to TCL[8], a state-of-the-art training-based method (Table 1). These results demonstrate the effectiveness of our loss function design in overcoming the missed and incorrect predictions from Stage 1, leading to significantly improved performance. In Appendix we show that both Stage 1.1 and Stage 1.2 are essential for training the alignment module.
Method Annotation mIoU
PV-21 PV-20 PC-59 CO-80 CS-171 City ADE
FMbSeg- Stage 1.1 - 25.37 42.92 22.03 32.97 16.96 9.50 10.45
FMbSeg- Stage 1.2 - 36.68 53.36 24.25 29.15 16.82 14.86 12.30
FMbSeg- Stage 1.1 (Semi-Supervised) Image-level 50.86 63.82 39.33 47.77 29.03 10.76 22.86
Table 2: Stage 1 ablation. Stage 1.2 achieves a better performance as it can segment small objects and multiple instances of the same object. Semi-supervised Stage 1.1 employs image-level labels for mask generation and achieves a comparable performance to SOTA training based methods.
4.4.2 Architecture Ablation
We design our alignment module as a single transformer block with multi-head self-attention over image patches. We ablate our choice against other designs, namely a single linear layer and a Multi-layer perceptron (MLP) with GELU activation. Table 3 reports the mIOU on CS-171. The differences are not substantial, with MLP achieving better performance than linear. The transformer block further improves over the MLP.
Architecture mIoU
Linear 27.25
MLP 28.90
Transformer block 29.88
Alignment Loss mIoU
$\ell_{SupCon}$ 26.25
$\ell_{\mathbf{t}}$ (2) 28.51
$\ell_{\text{TSupCon}}$ (1) 29.88
Table 3: CS-171 mIoU with our base model. Left. Comparison of different design choices for our alignment module, a transformer block has a small advantage. Right. Comparison of different losses. Our full loss TSupCon performs the best.
4.4.3 Loss Ablation
In section 3.3 we introduced our loss function for aligning patch features by contrasting them with each other and with text features. Here, we compare to two loss variants: 1) The supervised contrastive loss [20] (SupCon) alone where text features are considered as examples of the corresponding concepts similar to the image patches of a given category. 2) The prototype loss alone, Eq.2. Table 3 reports the mIOU on CS-171. We find that SupCon term alone exhibits the worst convergence while the prototype loss shows a stronger performance, probably due to the smooth approximation of Cross Entropy loss and SupCon loss on text-patch pairs. However, when combined with SupCon on pairs of patches, better performance is achieved due to more enhancement in the expressivity of these patches features. Overall, the unique treatment of TSupCon to the text features allows a smoother generalization, better convergence and hence considerably better performance.
4.5 Annotation-Free Segmentation Applications
To further illustrate the advantages of our open vocabulary segmentation tool, in this section we evaluate our annotation free semantic segmentation method on tasks different from those evaluated by common segmentation benchmarks. These serve as examples for different scenarios where it is required to segment a specific object with no pixel-level training data/pretrained model available.
4.5.1 Plug and Play Binary Segmentation Task
We consider a water segmentation task based on WaterV2 dataset from Kaggle [^1] https://www.kaggle.com/datasets/gvclsu/water-segmentation-dataset. Water’s uniform appearance, lighting conditions and anomalies (i.e. objects or reflections inside the water) make this a difficult segmentation task. Fig 5 shows some qualitative results for the water segmentation task. FMbSeg Stage 1 achieves an mIoU accuracy of $83.1\%$ on the evaluation set of this dataset.
4.5.2 LLM Personalization Task
Here we briefly showcase our method’s ability to fit into a completely different task. The main goal is to identify specific instances of objects such as personal items (My cat, my running shoes, my espresso cup etc.) given very limited number ($=4$) training views for each object. Figure 6 illustrates a few examples of the personalized objects in the training and evaluation set.
With no training, we employ our Stage 1 method to segment out each object in the training set given its name. For the test images, we query the model to generate a mask for each personalized objects on every image. we measure the cosine similarity of the DINOv2 features corresponding to the segmented mask for the personalized object with those extracted from the corresponding training images. A high similarity would indicate the presence of the object instance.
We compare our simple pipeline to a training based method where a new classification head is trained for each personalized object instances [1]. As seen in table 4, by just using a similarity threshold for detection, our stage 1 performs comparable to [1] on 29 personalized objects without any modification for this task.
Figure 5: Water segmentation results: Stage 1 accurately segments bodies of water in presence of anomalies and different lighting conditions achieving a $83.1\%$ mIoU accuracy.
Figure 6: Personalized object segmentation. Our Stage 1 method can segment out the personalized items with no training.
Method Precision Recall
MyVLM [1] 90% 96%
FMbSeg-Stage 1 91% 87%
Table 4: Personalized object retrieval. Our out of the box method performs comparable to MyVLM’s classification heads trained specifically for this task.
@@ 5 Conclusion and Future Work
Given the impressive performance of vision-language foundation models and their transferability to various downstream tasks, we consider the problem of open vocabulary annotation-free semantic segmentation. By composition of different foundation models, namely CLIP and SAM, we extract free pixel level annotations. We then propose a lightweight alignment module that projects the embedding of any arbitrary pretrained vision encoder to the text encoder space. Our method can be deployed as a plug-and-play customized alignment module for any semantic segmentation dataset with zero annotations. We show SOTA results demonstrate the effectiveness of foundational models compositionality. As future work, we want to investigate other image encoders and continuous fine-tuning on new categories. Additionally, we want to explore how our alignment module can improve VLM models that are based on CLIP to further strengthen their object localization capabilities.